### v4

# Two Types of Gluons in QCD: Re-interpretation of ALEPH and CMS Gluon Jet Data

Abstract
The Abelian decomposition of QCD tells that there are two types of gluons in QCD, the color neutral nurons and colored chromons, which have different color quantum numbers and thus behave differently. This strongly implies that QCD has two types of gluon jets, the nuron jet and chromon jet. One quarter of the gluon jets is made of the nuron jets which have sharper jet radius and smaller particle multiplicity, while three quarters of them are made of chromon jets which have broader jet radius and larger particle multiplicity. Moreover, the nuron jet has a distinct color flow which forms an ideal color dipole pattern, while the chromon jets have distorted dipole pattern. In this paper we provide five circumferential evidences of the existence of two types of gluon jets from the existing ALEPH and DELPHI data on $e\bar{e}\rightarrow Z\rightarrow b\bar{b}g$ three jet events and the CMS data on Pb-Pb heavy ion collision.
Abelian decomposition of QCD, restricted potential, valence potential, monopole potential, two types of gluons, nuron, chromon, nuron jet, chromon jet, anormaly in particle multiplicity of gluon jet, anormaly in energy fragmantation of gluon jets, correlation between the low particle multiplicity and large energy fragmantation in gluon jets, anormaly in the gluon quenching in heavy ion collision
pacs:
12.38.-t, 12.38.Aw, 11.15.-q, 11.15.Tk
I Introduction
With the asymptotic freedom and the subsequent experimental confirmation of the gluon jet, the quantum chromodynamics (QCD) has become widely accepted as the theory of the strong interaction [1, 2, 3]. But it has many unresolved problems. The color confinement problem known as the millennium problem is the most outstanding one. Two popular conjectures to resolve this problem are the monopole condensation [4, 5, 6] and Abelian dominance [5, 6, 7, 8]. The monopole condensation is based on the assumtion that, just as the Cooper pair (the electron pair condensation) confines the magnetic flux in superconductors, the monopole condensation (more precisely the monopole-antimonopole pair condensation) in QCD could confine the color electric flux. To prove this, however, we have to separate the monopole potential from the QCD gauge potential. Can we do this?
On the other hand the Abelian dominance assumes that only the Abelian part of the gluon potential can make the confinement. This is because the non-Abelian part must be confined since they are colored, so that they can not play the role of the confiner. This is just like the prisoner who should be jailed can not play the role of the jailer. To prove this, however, we should be able to separate the Abelian part of the QCD potential gauge independently. How can we do this?
There are other problems. Consider the proton made of three quarks. Obviously we need the gluons to bind the quarks in the proton. However, the quark model tells that the proton has no constituent gluon. If so, what is the difference between the binding gluon and constituent gluon? And how can we separate them gauge independently?
Moreover, the simple group theory tells that two of the eight gluons corresponding to the diagonal generators of the color gauge group must be color neutral, while the other six which correspond to the off-diagonal generators must carry the color. This strongly implies that there are two types of gluons, the color neutral ones and colored ones, which should behave differently. Can we distinguish them?
The bottom line here is to understand how different the non-Abelian dynamics of QCD is from the Abelian QED. Obviously all non-Abelian gauge groups have the maximal Abelian subgroups. Can we separate the Abelian subdynamics corresponding to the Maximal Abelian subgroup from the full non-Abelian dynamics? And how different is this Abelian subdynamics from QED? To answer this we should first be able to separate the Abelian subdynamics of QCD gauge independently. How can we do this?
The Abelian decomposition of QCD could resolve these problems nicely [5, 6, 9]. It decomposes the QCD gauge potential to two partsgauge independently, the restricted (Abelian) potential which nevertheless retains the full non-Abelian gauge degrees of freedom and the valence potential which transforms gauge covariantly. Moreover, it tells that the restricted potential is made of two parts, the non-topological Maxwellian Abelian potential and the topological Diracian monopole potential.
The Abelian decomposition does not change QCD but allows us to understand the color dynamics more clearly, by revealing the important hidden structures of the conventional QCD. It tells that QCD has the restricted QCD (RCD) made of the restricted potential which describes the Abelian subdynamics of QCD which nevertheless retains the full color gauge degrees of freedom. This tells that the Abelian subdynamics of QCD is different from QED. RCD is made of not only the non-topological Maxwellian potential but also the topological Diracian monopole potential.
Moreover, it allows us to interprete QCD as RCD which has the valence potential as the colored source. This has deep implications. First, this decomposes the QCD Feynman diagrams in such a way that the color conservation is explicit, decomposing the gluon propagators to the nuron and chrom propagators. Second, this implies that the nuron (like the photon in QED) plays the role of the binding gluon and the chromon (like the quark) plays the role of the constituent gluon.
Third, this tells that what makes QCD different from QED is the topological Diracian monopole potential, which strongly implies that the monopole is responsible for the color confinement. Finally, this puts QCD to the background field formalism which provides us an ideal platform to calculate the QCD effective potential and porve the monopole condensation [9, 10, 11]. This is because we can treat the restricted part as the classical background and the valence part as the fluctuating quantum field.
Figure 1: The Abelian decomposition of the QCD gauge potential. In (A) it is decomposed to the restricted potential (kinked line) and the chromon (straight line) potential. In (B) the restricted potential is further decomposed to the nuron (wiggly line) and the monopole (spiked line) potentials.
This means that, after the confinement sets in, the perturbative QCD is made of the non-topological Maxwellian Abelian potential which describes the color neutral gluon (the nuron) and the gauge covariant valence potential which describes the colored gluon (the chromon). In other words, there are two types of gluons in QCD, the color neutral nuron and colored chromon.
The Abelian decomposition is not just a theoretical proposition. The existence of two types of gluons can be verified by experiment [12, 13]. The purpose of this paper is to provide five circumferential evidences of the existence of two types of gluons, re-interpreting the existing ALEPH and DELPHI gluon jet data coming from the Z boson decay and CMS gluon jet data coming from heavy ion collision. We argue that the known gluon jet anormalies and puzzles in the ALEPH, DELPHI, and CMS gluon jet data are in fact the evidence of the existence of two types of gluons in QCD.
II Abelian decomposotion of QCD: A Brief Review
Consider the SU(2) QCD for simplicity
${\cal L}=-\frac{1}{4}\vec{F}_{\mu\nu}^{2}.$ (1)
Choosing an arbitrary direction $\hat{n}$ as the Abelian direction and selecting the potential which parallelizes the Abelian direction imposing the condition
$D_{\mu}\hat{n}=0,$ (2)
we can project out the restricted potential $\hat{A}_{\mu}$ which describes the Abelian subdynamics of QCD from the QCD potential [5, 6],
$\vec{A}_{\mu}\rightarrow\hat{A}_{\mu}=A_{\mu}\hat{n}-\frac{1}{g}\hat{n}\times\partial_{\mu}\hat{n}.$ (3)
Notice that the restricted potential is made of two parts, the non-toplogical Maxwellian potential and the topological Diracian monopole potential.
Figure 2: The Abelian decomposition of the gluon propagators in SU(3) QCD. In (A) and (B) the three-point and four-point gluon vertices are decomposed, and in (C) the quark-gluon vertices are decomposed, to the nuron and chromon propagators. Here the monopole does not appear in the diagrams because it describes a topological degree.
With this Abelian projection we recover the full potential $\vec{A}_{\mu}$ adding the valence potential $\vec{X}_{\mu}$ to $\hat{A}_{\mu}$,
$\vec{A}_{\mu}=\hat{A}_{\mu}+\vec{X}_{\mu},~~~(\hat{n}\cdot\vec{X}_{\mu}=0).$ (4)
This is the Abelian decomposition of $\vec{A}_{\mu}$. Notice that $\hat{A}_{\mu}$ retains the full non-Abelian gauge degrees of freedom and $\vec{X}_{\mu}$ transforms covariantly under the gauge transformation. The Abelian decomposition of the QCD gauge potential is graphically shown in Fig. 1.
The advantage of the Abelian decomposition is that it is gauge independent. Once the Abelian direction is chosen, it follows automatically. And we can choose any direction to be Abelian. The Abelian decomposition of the SU(3) QCD is a bit more compilcated, but straightforward. Here we simply show the Abelian decomposition of the gluon propagators of the SU(3) QCD graphically in Fig. 2.
The Abelian decomposition allows us to do many things that we cannot do in the conventional QCD. First, in the pertubative regime it allows us to decompose the QCD Feynman diagrams in such a way that color charge consevation is explicit [10, 11]. This is evident in Fig. 2. Second, it replaces the quark and gluon model of the hadron to the quark and chromon model [14, 15]. This is because the nuron plays the role of the binding gluon and the chromon plays the role of the constituent gluon. This clarifies the hadron spectroscopy, in particular the glueball spectrum, greatly.
Figure 3: The SU(3) lattice QCD calculation which establishes the monopole dominance in the confining force in Wilson loop. Here the confining forces shown in full, dashed, and dotted lines are obtained with the full potential, the Abelian potential, and the monopole potential, respectively.
Third, it tells that QCD has a non-trivial Abelian core, the restricted QCD (RCD) made of the restricted potential which describes the Abelian subdynamics of QCD but retains the full color gauge symmetry. This allows us to prove the Abelian dominance, that RCD is responsible for the confinement. Using the Abelian decomposition we can calculate the Wilson loop integral and show regourously that the restricted potential is responsible for the linear confing force [8].
Moreover, we can back up this with the lattice QCD calculation. Implementing the Abelian decomposition on the lattice, we can calculatethe Wilson loop integral numerically with the full potential, the Abelian potential, and the monopole potential separately, and show that they all generate the linear confining force [16, 17]. This tells that we only need the monopole potential to generate the confining force. The result of the lattice calculation is shown in Fig. 3. This confirms that the monopole is responsible for the color confinement.
Furthermore, it puts QCD to the background field formalism, because we can treat the restricted potential as the slow varying classical field and the valence potential as the fluctuating quantum field [9]. This provides us a natrual platform for us to calculate the QCD effective potential and demonstrate the monopole condensation. Treating RCD as the classical background and integrating out the valence potential under the monopole background, we can calculate the one-loop QCD effective potential and prove that the physical vacuum of QCD is given by the stable monopole condensation [10, 11]. The effective potential of the SU(3) QCD is shown in Fig. 4. This tells that in the perturbative regime (after the monopole condensation) QCD is described by two types of gluons, the nuron amd chromon.
Figure 4: The one-loop SU(3) QCD effective potential which shows that the true minimum of the potential is given by the Weyl symmetric monopole condensation.
But, of course, the most important point of the Abelian decomposition is that QCD in the perturbative regime has two types of gluons, the color neutral nurons and the colored chromons, which behave differently. If this is true, we must be able to show the existence of two types of gluons experimentally.
III Nuron Jet and Chromon Jet
The prediction of gluon in QCD and the experimental confirmation of the gluon jet was a milestone that justified QCD as the theory of the strong interaction [2, 3]. This implies that, if there really exist two types of gluons, QCD must have two types of gluon jets.
Experimentally the gluon jet tagging has been a very complicated business which is not completely understood yet [18, 19, 20, 21, 22, 23, 24]. The success rate of the gluon jet tagging has been only about 70%. Moreover, the existing gluon jet data are known to have “anormalies” and “puzzles” which could not be explained within the framework of the conventional QCD [25, 26, 27]. In the following we argue that these anormalies and puzzles in the gluon jet experiments could be explained by the existence of two types of gluons. In particular, we provide five circumferential evidences which strongly imply the existence of two types of gluons from the existing ALEPH, DELPHI, and CMS gluon jet data.
To confirm the existence of two types of gluon jets experimentally, we have to know the basic features of the nuron and chromon jets. Since two of the eight gluons are nurons, one quarter of the gluon jets should be the nuron jets and three quarters of them should be the chromon jets. If so, how can we test the existence of two types of gluon jets experimentally?
To answer this remember that the gluons and quarks emitted in the p-p collisions evolve into hadron jets in two steps, the soft gluon radiation of the hard partons described by the perturbative process and the hadronization described by the non-perturbative process. Since the hadronization in the second step is basically similar in all jets, the main differences come from the parton shower (the soft gluon radiation) in the first step. From the Abelian decomposition of the Feynman diagram shown in Fig. 2 we can figure out the difference of the soft gluon radiation of the nuron, chromon, and quark jets. This is shown in Fig. 5.
Figure 5: The parton shower (the soft gluon radiation) of hard partons. The strength of the parton shower of the nuron jet shown in (A) is of $O(g^{2})$, but the chromon jet and the quark jet shown in (B) and (C) are of $O(g)$. Moreover, the nuron jet has only one type of the parton shower, while the chromon and quark jets have five and two.
Clearly the nuron jet shown in (A) has only one type of soft gluon radiation (i.e., the two-chromon radiation). In contrast, the chromon jet in (B) has five different types of soft gluon radiations, and the quark jet in (C) has two. Moreover, the leading order of the soft gluon radiation of the nuron jet shown in (A) is of $O(g^{2})$, while that of the chromon and quark jets shown in (B) and (C) are of the order of $O(g)$.
This tells that the nuron jet is expected to behave more like the photon jet, sharper than the chromon and quark jets with relatively smaller jet radius. This is because the nuron has only one soft gluon radiation, while the chromon has four more which could only broaden the jet. Moreover, this also strongly implies that the nuron jet should have considerably less particle multiplicity than the chromon and quark jets. This must be clear from Fig. 5.
The soft gluon radiation of the nuron jet shown in (A) has another important implication. Since it has only one type of parton shower made of chromon-antichromon pair, it could most likely produce glueball states (more precisely the chromoball states) made of the chromon-antichromon pair, if the energy of the jet becomes close to the mass of the chromoball states [14, 15]. On the other hand, the chromon and quark jets shown in (B) and (C) imply that they have little chance to produce such states. This implies that (when the jet energy becomes close to the glueball mass) the nuron jets could produce more neutral particles than the chromon and/or quark jets, because the chromoballs produce more neutral particles [14, 15]. This provides another characteristic feature of the nuron jet. As we will see, this point will have an important consequence in the following.
Figure 6: The expected gluon jet distribution against the particle multiplicity. Here we have plotted the overall particle multiplicity when the distance between the chromon and nuron peaks becomes one, two, and three standard deviations. Notice that the black curve represents the well known particle multiplicity we have in the conventional QCD.
Suppose the nuron jet has less particle multiplicity. In this case there should be one more local peak on the side of the less number of particles in the particle distribution in the gluon jet. So, the particle multiplicity of the gluon jets must have an asymmetry against the main peak axis toward the less number of particles [13]. The expected particle multiplicity distribution of the gluon jets is shown in Fig. 6. In this figure we have assumed (for simplicity) that the particle distribution of nuron and chromon jets are Gaussian, and plotted the expected particle distribution when the nuron peak is one, two, and three standard deviations away from the chromon peak. Notice that the shape of the overall particle distribution of the gluon jets crucially depends on the distance between the two peaks. But clearly the figure shows that we have more particles in the left side of the peak.
For example, when the nuron peak is three standard deviation away, the left part of the peak axis has 64 % more particles than the right part. And we have 42 % more events when the nuron peak is two standard deviation away. But when the distance becomes one standard deviation, the left part has only 6 % more particles. So the nuron jet could easily be left unnoticed when the distance between the nuron and chromon peaks becomes less than two standard deviation.
This should be contrasted with the black curve, which represents the particle distribution of the gluon jets when the nuron and chromon jets have the same particle multiplicity. This, of course, is the well known particle distribution of the gluon jets we expect in the conventional QCD (without the Abelian decomposition). In this case the asymmetry disappears completely. This tells that the existence of the reflection asymmetry (tilt) of the particle distribution against the peak axis, if exists, can be viewed as a strong indication that there could be two types of gluon jets.
Figure 7: The ALEPH data on the energy fragmantation of particles of the gluon and quark jets. The Monte Carlo predictions of Jetset and Herwig are shown in solid and dotted lines, and the experimental data of the gluon and quark jets are shown in square and solid points. Clearly the data shows an unmistakable excess of particles with largy energy fragmatation of gluon jets.
There is another way to confirm the existence of the nuron jets. Notice that the equipartition of energy tells that nuron jets and chromon jets must have the same energy in average. So, if the nuron jets have less particle multiplicity as we expect, the particles coming from the nuron jets must have more energy in geneal than those coming from the chromon jets. If so, checking the energy distribution of the particles coming from the gluon jets and verifying the existence of the excess of particles with higher energy, we can show the existence of the nuron jets.
This means that, without trying to prove the existence of the nuron jet and the chromon jet directly, we could tell the existence of two types of gluon jets, checking the particle multiplicity and the energy distribution of the particles from the existing gluon jets data [12, 13]. This is really remarkable.
IV Experimental Evidence of Two Types of Gluons: Re-interpretation of ALEPH, DELPHI, and CMS Gluon Jet Data
The above discussion tells that one quarter of the existing gluon jets should actually be the nuron jets which have different jet shape (the sphericity), particle multiplicity, and color dipole pattern. Is there any evidence for this in the existing gluon jet data? The answer is definitely yes.
Figure 8: The leftside is the ALEPH data on the charged particle multiplicity of the gluon jets. The Monte Carlo prediction of Jetset is shown in solid line and the experimental data of the gluon jets are shown in solid points. The rightside is the integral of the events of the ALEPH data, which shows an asymmetry against the peak axis at $n_{c}=4$.
A best place to look for such evidence is the ALEPH gluon jet data coming from the $Z$ boson decay [25]
$e\bar{e}\rightarrow Z\rightarrow b\bar{b}g.$ (5)
The ALEPH collaboration has the published results which indicate the existence of two types of gluons [25], the “anormaly” in the energy fragmantation of the particles shwn in Fig. 7 and the “abnormal” charged particle multiplicity of the gluon jets shown in Fig. 8. Consider the energy fragmantation of the particles first. Clearly Fig. 7 shows that the Monte Carlo predictions of the energy distribution of the particles coming from the quark jet are in good agreement with the experimental data, but those of the gluon jets have an obvioius disagreement with the experiment, the unexpected excess of the more energetic particles shown in red circles. This is the well known anormaly of the energy fragmantation of the particles in the gluon jets.
But as we have pointed out, this anormaly is in fact exactly what we would expect if we have two types of gluons. This is because the particles from the nuron jets should have more energy than those from the chromon jets, since the nuron jets produces less particles. This is precisely what we observe in the ALEPH data in Fig. 7. This strongly implies that the anormaly is not an anormaly but an evidence of the existence of the nuron jet.
The same ALEPH data provides us another evidence of two types of gluon jets, on the (charged) particle multiplicity shown in the leftside of Fig. 8. According to our prediction this ALEPH data must have the reflection asymmetry against the peak axis, the excess of the less particles, predicted in Fig. 6. Remarkably, the ALEPH data does show this asymmetry. It tells that the peak is located at $n_{C}=4$, but clearly shows the excess of the events against the Monte Carlo prediction when $n_{c}$ becomes less than 4.
Figure 9: The DELPHI experiment which shows a clear excess of the neutral particles in gluon jet in (a) and (c), which is absent in the quark jet in (b) and (d).
Moreover, the integral of the experimental data shown in the rightside of the figure tells that the particle multiplicity is asymmetric (tilted leftward) against the peak axis. It shows that 42 % of the total number of the charged particles are less than 4 but 22 % are more than 4. This tells that the leftside of the peak has 20% more jets. In fact, according to our analysis shown in Fig. 6, the asymmetry implies that the experimental data could be made of two peaks, the chromon peak with 75 % population near $n_{C}=4$ and the nuron peak with 25 % population located leftside roughly 1.5 standard deviation away from the chromon peak.
Suppose ALEPH discovered this nuron peak without knowing the existence of the nuron jet. In this case they would have called this an “anormaly”, because they would have no way to explain this. So we could say that the old ALEPH gluon data actually have two anormalies, the anormaly in energy fragmantation and the anormaly in particle multiplicity. But this second anormaly (in particle multiplicity) has been completely unnoticed till now. And this new anormaly in the particle multiplicity of the ALEPH gluon jets could be viewed as another evidence of the nuron jet.
So far we have discussed the gluon jet anormalies of the ALEPH data. But there exists another type of the gluon jet anormaly coming from the DELPHI experiment [26]. Based on the same $Z$ boson decay shown in (5), the DELPHI group separated the quark jets with purity about 90 % and the gluon enriched jets with purity about 70 %, and observed a clear excess of the neutral particles in the gluon jets at low invariant masses less or equal to 2 GeV. This is shown in Fig. 9. Obviously this is an anormaly. The DELPHI group has interpreted this anormaly as “an indication that the gluon jet might have an additional hithertoo undetected fragmantation mode via a two-gluon system” telling that “this could be an indication of a possible production of gluonic states as predicted by QCD” [26].
Figure 10: The expected gluon quenching in heavy ion collision. The gluons in p-p collision can easily escape because the volume of the quark-gluon plazma is small, but in Pb-Pb collision they are supposed to be quenched before they escape. Notice, however, the nurons can easily escape because the coupling is weak.
However, this is precisely the characteristic feature of the nuron jet that we have mentioned above. Obviously the neuron jet shown in (A) of Fig. 5 tells that (when the jet energy is close to the glueball mass) it could most likely produce the glueballs (more precisely, the chromoballs) made of the chromon-antichromon pair, which produce mostly the neutral particles. This shows that the figure in (A) of Fig. 5 is not only in line with the DELPHI interpretation, but more importantly provides a theoretical justification of this interpretation. Moreover, this tells that the excess of the neutral particles in the gluon jet comes from the existence of the neuron jet, which could be interpreted as one more supporting evidence of the two types of gluon jets. This is remarkable.
We have another circumferential evidence for two types of gluons coming from an unexpected quarter, the CMS data on the heavy ion collision [27]. The gluon jets in heavy ion collisions in general are supposed to undergo the quenching when they pass through the quark-gluon plasma of the heavy ions, so that the gluon jets coming out from the heavy ion collisions are expected to be less than those coming from the p-p collision. This is the quenching effect schematically shown in Fig. 10. But in the CMS experiment of the Pb-Pb collision, the gluon jets coming from the Pb-Pb and p-p collisions have not much difference, which implies that the gluons undergo less quenching than expected in heavy ion collision. The CMS result is shown in Fig. 11. This has been thought to be a “puzzle” [27].
Fortunately, this unexpected puzzle could also be explained in the framework of the Abelian decomposition. Since the nurons are color neutral their interaction with the quark-gluon plasma must be weaker compared to the chromon interaction, as we have explained in Fig. 5. So the nurons come out almost unquenched in the Pb-Pb heavy ion collision. This is shown in Fig. 10. And this could explain the puzzle in the CMS heavy ion collision observed in Fig. 11.
So far we have discussed four circumferential evidences for the existence of two types of gluon jets. Each one separately may not be viewed as serious, but together they do make a serious case. One could go further and check if the ALEPH data has two jet shapes (sharper nuron jets and broader chromon jets) and different color dipole patterns (ideal color dipole pattern for the nuron jets and distorted dipole pattern for the chromon jets).
Figure 11: The CMS gluon jet data in Pb-Pb heavy ion collision compared to the gluon jets in p-p collision. The expected gluon quenching in Pb-Pb collision is hardly detectable.
But in our mind a more important test is to check the correlation between the two ALEPH data shown in Fig. 7 and Fig. 8. Notice that the data which show the anormaly in Fig. 7 and the asymetry in Fig. 8 actually come from the same nuron jets, so that they should be thought as different aspects of the same physics. This means that, if our interpretation is correct, they should be correlated. So, confirming this correlation we could increase the credibility of the above interpretation greatly. Unfortunately, ALEPH had no result on this correlation, probably because they had no motivation to check this correlation.
V Re-analysis of ALEPH Gluon Jet Data by MIT-UoS Group
The above ALEPF, DELPHI, and CMS gluon jet data stongly indicate the existence of two types of gluons. On the other hand, these anormalies have not been confirmed by an independent group, so that the data need to be re-analyzed. Recently the MIT-UoS group has reanalyzed the ALEPH data. The purpose was two-fold: Is there really an anormaly in the gluon jet data? And if so, does this imply the existance of two types of gluons? Now, we discuss the outcome of the re-analysis [28].
Figure 12: The MIT-UoS reanalysis of the ALEPH gluon jet data on the total number of the particle multiplicity. The result confirms the excess of the gluon jet events for $n=1,2$, and 333.
From the 2900317 raw ALEPH $b\bar{b}g$ three jet events coming from the $Z$ boson decay obtained from the years 1991 to 1995, the MIT-UoS group selected 40984 gluon jets using the gluon tagging method similar to the one adopted by the ALEPH. With this they obtained the particle multiplicity distribution of the gluon jets for all, charged, and neutral particles. The result of the particle multiplicity for all particles is shown in Fig. 12. It has the peak at $N=6$, but also shows that the number of the events around $N=2$ and 3 is roughly 2 and 1.5 times more than the Monte Carlo prediction. This strongly indicates the existence of a second peak around $N\simeq 2.5$. This is precisely the anormaly of the particle multiplicity distribution coming from the nuron jet that we have discussed.
They also have re-analyzed the charged and neutral particle multiplicity distribution, and confirmed the existence of the same trend. For example, for the charged particle multiplicity distribution, the peak is located at $N=4$, but it has a second peak at $N=1$ which indicates the existence of the anormaly in the charged particle multiplicity distribution. With this we could conclude that the anormaly in the particle multiplicity of the ALEPH gluon jet data is real, and that we could attribute this anormaly to the existence of the nuron jets. It is really interesting that the MIT-UoS result is virtually identical to our re-interpretation of the published result of the ALEPH jet data shown in Fig. 8.
With the same ALEPH data, the MIT-UoS group has re-analized the energy fragmantation of all, charged, and neutral particles coming from the gluon jets. The result of the energy fragmantation of all particles is shown in Fig. 13, which should be compared with the ALEPH result shown in Fig. 7. Here again, the result shows that the gluon jet energy fragmantation above 0.5 is 1.5 to 2 times bigger than the Monte Carlo prediction, which confirms the existence of the anormaly of the ALEPH gluon jets in large $X_{E}$ above 0.5. And obviously, this could be attributed to the existence of the nuron jets which have more energetic particles.
Figure 13: The MIT-UoS re-analysis of the ALEPH gluon jet data on the distribution of the energy fragmantation of the particles. The result confirms the excess of the energy fragmantation $X_{E}$ above 0.5.
If this interpretation is correct, the two anormalies should be correlated, because they come from the same nuron jets. Remarkably, we can demonstrate this correlation from the ALEPH data. The MIT-UoS group have checked this correlation between all, charged, and neutral particle multiplicity and the corresponding energy fragmantation, and confirmed the existence of the correlation. The correlation for all particles is shown in Fig. 14.
It shows that 32 % of the events below $N=6$ fully expain the anormaly in the large energy fragmantation. Indeed, since this 32 % also contains the tails of the chromon jets which have high particle multiplicity, we can tells that the nuron jets which consist of 25 % of the gluon jets with the low particle multiplicity less than $N=6$ can fully account for the large energy fragmantation over 0.5 shown in Fig. 13. This confirms that there exists a strong correlation between the low particle multiplicity events and high energy fragmantation events. This is precisely the correlation predicted by the existence of the nuron jets. The MIT-UoS group also confirmed the existence of the similar correlation for the charged and neutral particles.
Clearly this confirmation of the correlation between the low particle multiplicity events and high energy fragmantation events is the most compelling evidence of the exisence of the nuron jet. Of course, the excess of the low particle multiplicity shown in Fig. 8 and Fig. 12 or the excess of the high energy fragmantation shown in Fig. 7 and Fig. 13 independently implies the existence of the nuron jets. However, there is no reason why they should be correlated, unless they come from the same origin as we have predicted based on the Abelian decomposition of QCD. And the above correlation assures that indeed our interpretation is correct.
Figure 14: The correlation of the gluon jets between the charged particle multiplicity and the energy fragmantation of particles of the ALEPH experiment.
VI Discussion
The Abelian decomposition of QCD tells that there are two types of gluons, the color neutral nurons and colored chromons, in QCD. In this paper we have studied the four gluon jet anormalies and puzzles in the ALEPH, DELPHI, and CMS gluon jet experiments, and argued that these anormalies and puzzles could in fact be interpreted as the evidence of the existence of two types of the gluon jets. Based on the Abelian decomposition of QCD, we have argued that the excess of the low particle multiplicity and high energy fragmantation in the ALEPH data, the excess of the neutral particles in the gluon jets in the DELPHI data, and the absence of the gluon quenching in the Pb-Pb collision in the CMS data could be interpreted as the circumferential evidence of two types of gluons in QCD.
More importantly, we have discussed the existence of a new correlation between the low particle multiplicity events and high energy fragmantation events in the ALEPH data, which is obtained by the MIT-UoS group re-analyzing the ALEPH gluon jet data [28]. This correlation can be viewed a most important and persuasive evidence of the existence of the nuron jet. This is because this shows that both excesses come from the same origin, the nuron jet.
Admittedly, the above evidences may not be viewed as a direct evidence of the existence of two types of gluons, since they do not identify the nuron and chromon jets separately. To have the direct evidence, we need new Monte Carlo simulations, new Pythia and FastJet predictions based on the Abelian decomposition of QCD, and confirm the existence of two types of gluon jets. This is beyond the scope of this paper. Nevertheless the above circumferential evidences strongly implies that indeed there could be two types of gluons in QCD. This provides us an irresistable motivation to re-analyse all existing gluon jet data more carefully.
There are different ways to test the existence of two types of gluon jets. One is the jet shape. As we have pointed out, the nuron jet is expected to have a sharper jet shape, while the chromon jet has broader jet shape. So we could test the existence of the nuron jet studying the jet shape of the gluon jets. Another important feature of the nuron jet is that it has different color flow. Clearly the chromons and quarks carry color charge, but the nurons are color neutral. So the nuron jet must have different color flow. In fact, Fig. 5 suggests that the nuron jet generates an ideal color dipole pattern, but the other two jets have distorted dipole pattern. But as far as we understand, these studies have not been done by any experimental group.
It should be emphasized that, in doing so we do not need any new experiment. LHC produces billions of hadron jets in a second, and ATLAS and CMS have already filed up huge data on gluon jets. Moreover, DESY, LEP, and Tevatron have old data on three jet events which could be used to confirm the existence of the nuron and chromon jets. Here again the simple number counting strongly suggests that one quarter of the gluon jets coming from the three jets events should actually be the nuron jets which do not fit to the conventional gluon jet category.
There have been many experiments on the gluon jets, and one might wonder why they could not find two types of the gluon jets. There could be many reasons, but the most important reason might be that there was no theoretical motivation for them to do so, probably because they were unaware of the possibility of two types of gluons in QCD.
As we have emphasized, the Abelian decomposition does not change QCD, but reveals the hidden structures of QCD. It tells that QCD has an Abelian core RCD, and QCD can be viewed as RCD which has the chromons as colored source [5, 6]. It decomposes the QCD Feynman diagrams in such a way that color conservation is explicit [9]. It allows us to prove the Abelian dominance and the monopole condensation [10, 11]. And it replaces the quark and gluon model by the quark and chromon model [14, 15].
But most importantly, it tells that there are two types of gluons, the color neutral nurons and colored chromons which behave differently. A common prejudice in the conventional QCD is that the eight gluons are supposed to be indistinguishable, and nobody has doubted this popular wisdom. But the Abelian decomposition tells that this may not be true. In this paper we challenged this wisdom and showed how to back up this experimentally by discussing five evidences of two types of gluons. Our findings in this paper strongly implies that we may need a totally new interpretation of QCD.
At this point one might ask what are the gluon jets identified so far by experiments. Most probably they are the chromon jets, because the chromon jet has the characteristics of the gluon jet in the conventional QCD. This is evident from Fig. 5. A more interesting question is why they have not found the nuron jet. Obviously, they have not seriously searched for the nuron jet, because they had no motivation to do so. Moreover, the nuron jets consists of only 25 % of the gluon jets, and might have been systematically excluded because they do not fit to the typical gluon jet.
In fact, the sucess rate of the gluon jet tagging have been only around 70 %, which implies that the experiments have about 30% of uncertainty on the gluon jet. We believe that one reason for this uncertainty is because 25 % of the gluon jets are actually made of the nuron jets which behave differently. So it is really remarkable that the ALEPH, DELPHI, and CMS gluon jet data were able to record evidences for two types of gluon jets in spite of this uncertainty.
In this paper we have discussed the experimental verification of the Abelian decomposition. But the physical applications of the Abelian decomposition is as important as the experimental verification. Obviously a natural place to apply the Abelian decomposition is the parton distribution in QCD and have a new parton distribution function for hadrons, in particular for the proton parton distribution function. In this direction there has been an important progress. Recently the theory group of Institute of Modern Physics of Chinese Academy of Science has re-analized the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution equation of the gluon parton distribution in proton, and suceeded to decompose it and obtain two new evolution equations for the nuron and chromon saparately [29]. This could change our understanding of the parton distribution and evolution in QCD drastically. For this to happen, of course, we first need to confirm the existence of two types of gluons in QCD experimentally. And this was precisely what we tried to do in this paper.
The experimental discovery of the gluon jet was a big step for us to understand QCD [1, 2]. Clearly the confirmation of two types of gluons will be another giant step for QCD. If confirmed, this could be a most important discovery in QCD which could extend the horizon of our understanding of nature one step further.
Acknowledgements
The author thanks Yen-Jie Lee for informing the CMS data on the gluon quenching anormaly in heavy ion collision, and Inkyu Park and Youngkwon Jo for sharing the result of the MIT-UoS re-analysis of the ALEPH data and allowing to use some of them in this paper. The work is supported in part by National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University.